# Fokker-Planck approach for thermal fluctuations in antiferromagnetic systems

### v1

@@ I Introduction
Two-dimensional (2D) magnetic materials have recently emerged as a frontier in condensed matter physics, offering a unique platform for exploring fundamental spin physics and developing next-generation spintronics du2020berry; zhang2020gate; wang2023intrinsic; pancholi2025journey; liu2020robust. The experimental ability to isolate atomically thin layers in van der Waals crystals has enabled direct access to intrinsic magnetic order in materials where the reduced dimensionality and enhanced fluctuations reveal novel magnetic phenomena ghosh2021low; liang2023small; wu2023delaminated; ghosh2024electronic. Among these, 2D antiferromagnets (AFMs) are of particular interest due to their unique properties, such as ultrafast dynamics, minimal stray fields, and robustness against external magnetic fields, which makes them attractive for information storage and processing baltz2018antiferromagnetic; gomonay2014spintronics. In particular, transition-metal phosphotrichalcogenides, denoted MPX3 (usually with M = Fe, Ni, Mn and X = S, Se) are van der Waals-layered AFM semiconductors with interesting magnetic and spintronic properties mayorga2017layered; wang2018new; samal2021two; ghosh2021low. Understanding the spin dynamics and the role of thermal fluctuations in these low-dimensional systems is crucial, as their low dimensionality enhances thermal fluctuation. At the same time, their intrinsic uniaxial anisotropy stabilizes the AFM order.
@@
The dynamics of magnetization in magnetic systems is typically described by the Landau-Lifshitz-Gilbert (LLG) equation, which accounts for precession and damping of the magnetization under the action of an external magnetic field gilbert1955lagrangian; gilbert2004phenomenological; kamra_prb_2018; brown_prev_1963. In AFM materials, the structure made of two sublattices leads to coupled LLG equations that in turn give rise to interesting physical effects, e.g. mode-dependent damping and spin-pumping liu2017mode; kamra_prb_2018. At finite temperature, one can extend the LLG equation to account for thermal fluctuations by incorporating Langevin random fields into the LLG equation, producing a stochastic LLG equation garanin1997fokker; chotorlishvili_prb_2013; atxitia_prb_2012.
@@
In this work, starting from a stochastic LLG equation, we develop and apply a comprehensive Fokker-Planck (FP) formalism tailored for describing fluctuations and dynamics in an AFM system with uniaxial anisotropy. Our primary goal is to use the FP equation, coupled with a mean-field (MF) approximation, to derive the equations of motion for the spin polarization and the two-time spin-spin correlation functions. This approach allows us to incorporate thermal fluctuation corrections.
@@
The results of this formalism are exploited to analyze the spin-wave dynamics, providing a renormalization of characteristic energies and damping rates induced by thermal fluctuations in the AFM phase. Moreover, we have developed a phenomenological model to describe the influence of magnetic fluctuations on the resistance noise of 2D antiferromagnetic semiconductors, motivated by recently observed anomalous resistance fluctuations in MPX3 near the Néel temperature ghosh2021low; ghosh2024electronic.
@@ II Model
In this Section, we first present a classical model for describing a simple AFM with uniaxial anisotropy rezende2019introduction. We then introduce a stochastic Landau-Lifshitz equation derived from this model, in which Langevin fields are incorporated to account for thermal fluctuations. This stochastic approach leads to a FP formalism that describes the staggered magnetization dynamics under the influence of thermal fluctuations.
@@ II.1 Magnetism
We introduce the model to describe an AFM composed of two interpenetrating classical spin sublattices (A,BA,Bchen2019landau; rezende2019introduction, whose magnetic properties are described by the Hamiltonian
$\mathcal{H}$ $=-\sum_{\alpha=A,B}\sum_{p}\bm{h}^{\rm ext}_{{\alpha p}}\cdot{\bf\it s}_{\alpha p}+\mathcal{J}\sum_{\expectationvalue{p,q}}{\bf\it s}_{Ap}\cdot{\bf\it s}_{Bq}$ (1)
$+\mathcal{D}\sum_{\expectationvalue{p,q}}s_{Ap,z}s_{Bq,z}~,$
where the first term is the Zeeman term, representing the interaction with an external magnetic field $\bm{h}^{\rm ext}$, the second term characterizes the exchange interaction energy, and the final term considers the possible anisotropy present in the spin-spin interaction. In this study, we focus exclusively on uniform spin-spin interactions with strength $\mathcal{J}>0$, between nearest-neighbour sites, denoted as $\expectationvalue{p,q}$. The unit vector ${\bf\it s}_{\alpha p}$ denotes the orientation of the spin located on the site $p$ on the sublattice $\alpha\in\{A,B\}$. The parameter $\mathcal{D}$ takes into account the fact that the system can have a cylindrical symmetry due to the geometry of the configuration, and represents the interaction anisotropy between the different sublattices, attributed to their antiparallel configuration. In particular, for $\mathcal{D}>0$, this model favors alignment along the $z$ direction, sometimes referred to axial anisotropy. Conversely, for $\mathcal{D}<0$ one obtains a stronger interaction in the $x$-$y$ plane. The spherical symmetry is recovered by taking $\mathcal{D}=0$. The effective field acting on a given spin at the site $p$ of the sublattice $\alpha$ is expressed as
$\bm{h}_{\alpha p}=-\frac{\partial\cal H}{\partial\bm{s}_{\alpha p}}=\bm{h}^{\rm ext}_{\alpha p}-{\cal J}\sum\nolimits^{\prime}_{q}\bm{s}_{\bar{\alpha}q}-{\cal D}\sum\nolimits^{\prime}_{q}s_{\bar{\alpha}q,z}\bm{u}_{z}~,$ (2)
where $\bar{\alpha}=B$ ($\bar{\alpha}=A$) for $\alpha=A$ ($\alpha=B$), $\bm{u}_{z}$ is the unit vector directed along the $z$-direction, $\sum\nolimits^{\prime}_{q}$ represents the summation over the sites $q$ of the sublattice $\bar{\alpha}$ that are nearest neighbours to the site $p$ of the sublattice $\alpha$.If the exchange interaction dominates over the anisotropies and external fields, the spatial variations are smooth, and the thermal fluctuations are small, we can apply the macrospin approximation rezende2019introduction. For each sublattice, we have a collective magnetization, ${\bf\it M}_{\alpha}=N_{\alpha}{\bf\it s}_{\alpha}$, where $N_{\alpha}$ being the spin density on the sublattice $\alpha\in\{A,B\}$, and for simplicity we assume $N_{A}=N_{B}=N$. When disregarding the external magnetic field and applying the macrospin approximation, the effective field assumes a homogeneous form and can be rewritten as
@@
with $J=Z\cal J$ and $\Delta=Z\cal D$, where $Z$ denotes the coordination number that considers the number of nearest neighbors for a generic site.
@@ II.2 Fokker-Planck formalism
Here, we derive the dynamics equation for the probability distribution of the magnetization configuration of an AFM, starting from a stochastic version of the LLG equation that includes Langevin fields to incorporate the effects of thermal fluctuations  garanin1997fokker. For each of the two sublattices constituting the AFM, the stochastic LLG equations are expressed as
@@
where $\dot{{\bf\it s}}_{\alpha}=\partialderivative{{\bf\it s}_{\alpha}}{t}$, $\alpha\in\{A,B\}$, $\bm{h}_{\alpha p}$ is the effective magnetic field associated with the Hamiltonian of the spin defined in Eq. (2), $\bm{\zeta}_{\alpha p}$ is a Langevin field that is postulated to be uncorrelated at different sites. Note that we are working in units such that the lattice-independent chen2019landau, gyromagnetic ratio $\mu=1$, and $\hbar=1$. Both $\Lambda$ and $\Lambda^{\prime}$ are two diagonal matrices
@@
where $\lambda,\lambda+\delta\lambda\ll 1$ are the intra-sublattice Gilbert damping factors gilbert1955lagrangian; gilbert2004phenomenological, and $\lambda^{\prime},\lambda^{\prime}+\delta\lambda^{\prime}\ll 1$ are the inter-sublattice Gilbert damping factors kamra_prb_2018. Inclusion of parameters $\delta\lambda$ and $\delta\lambda^{\prime}$ allows uniaxial symmetry in damping mechanisms, following the cylindrical symmetry of the spin-spin interactions.
@@
Following Ref. garcia2007introduction, Eq. (6) can be rewritten in the form of the (componentwise) Langevin equation
@@
Langevin fields $\bm{\zeta}_{\alpha}$ are considered stochastic processes titov_jmm_2021; garcia2007introduction; garanin1997fokker. Assuming that they are Gaussian and that they satisfy the following statistical properties
@@
Moreover, we underline that the condition $D_{\mathrm{T},j}\geq 0$ leads to a constraint on the anisotropy $\delta D_{\mathrm{T}}\geq-D_{\mathrm{T}}$. Using the quantities in Eqs. (8) and (9), it is possible to obtain the coefficients of the Kramers-Moyal expansion, which describes the evolution of the probability distribution function (PDF) for a generic Markov process kramers1940brownian; moyal1949quantum; garcia2007introduction (see Appendix A for further details). One obtains the FP equation garcia2007introduction, namely the equation governing the dynamics for the PDF, $\mathcal{P}({\bf\it\mathcal{S}},t)$,

### v2

@@ I Introduction
Two-dimensional (2D) magnetic materials have recently emerged as a frontier in condensed matter physics, offering a unique platform for exploring fundamental spin physics and developing next-generation spintronics [9, 52, 48, 39, 31]. The experimental ability to isolate atomically thin layers in van der Waals crystals has enabled direct access to intrinsic magnetic order in materials where the reduced dimensionality and enhanced fluctuations reveal novel magnetic phenomena [14, 29, 50, 15]. Among these, 2D antiferromagnets (AFMs) are of particular interest due to their unique properties, such as ultrafast dynamics, minimal stray fields, and robustness against external magnetic fields, which makes them attractive for information storage and processing [2, 19]. In particular, transition-metal phosphotrichalcogenides, denoted MPX3 (usually with M = Fe, Ni, Mn and X = S, Se) are van der Waals-layered AFM semiconductors with interesting magnetic and spintronic properties [34, 47, 45, 14]. Understanding the spin dynamics and the role of thermal fluctuations in these low-dimensional systems is crucial, as their low dimensionality enhances thermal fluctuation. At the same time, their intrinsic uniaxial anisotropy stabilizes the AFM order.
@@
The dynamics of magnetization in magnetic systems is typically described by the Landau-Lifshitz-Gilbert (LLG) equation, which accounts for precession and damping of the magnetization under the action of an external magnetic field [16, 17, 24, 3]. In AFM materials, the structure made of two sublattices leads to coupled LLG equations that in turn give rise to interesting physical effects, e.g. mode-dependent damping and spin-pumping [32, 24]. At finite temperature, one can extend the LLG equation to account for thermal fluctuations by incorporating Langevin random fields into the LLG equation, producing a stochastic LLG equation [11, 6, 1].
@@
In this work, starting from a stochastic LLG equation, we develop and apply a comprehensive Fokker-Planck (FP) formalism tailored for describing fluctuations and dynamics in an AFM system with uniaxial anisotropy. Our primary goal is to use the FP equation, in a mean-field (MF) approximation, to derive the equations of motion for the spin polarization and the two-time spin-spin correlation functions. This approach allows us to incorporate thermal fluctuation corrections.
@@
The results of this formalism are exploited to analyze the spin-wave dynamics, providing a renormalization of characteristic energies and damping rates induced by thermal fluctuations in the AFM phase. Moreover, we have developed a phenomenological model to describe the influence of magnetic fluctuations on the resistance noise of 2D antiferromagnetic semiconductors, motivated by recently observed anomalous resistance fluctuations in MPX3 near the Néel temperature [14, 15].
@@ II Model
In this Section, we first present a classical model for describing a simple AFM with uniaxial anisotropy [44]. We then introduce a stochastic Landau-Lifshitz equation derived from this model, in which Langevin fields are incorporated to account for thermal fluctuations. This stochastic approach leads to a FP formalism that describes the staggered magnetization dynamics under the influence of thermal fluctuations.
@@ II.1 Magnetism
[…]
Figure 1: Schematics of the three antiferromagnetic ordering patterns in a single-layer honeycomb lattice: a) Néel, b) zigzag, and c) stripy configurations. Orange and blue circles denote opposite spin orientations. The purple lines indicate the nearest (dotted line), second-nearest (dash-dotted line), and the third-nearest-neighbor sites (dashed line) of the central site.
We model the magnetic properties of a two-dimensional antiferromagnet by an effective classical spin Hamiltonian defined on a lattice [5],
${\cal H}=$ $-\sum_{j}\bm{h}_{j}^{\mathrm{ext}}\cdot\bm{s}_{j}+\sum_{\langle j,k\rangle}\left(\mathcal{J}_{1}\bm{s}_{j}\cdot\bm{s}_{k}+\mathcal{D}_{1}s_{j,z}s_{k,z}\right)$ (1)
$+\sum_{\langle\langle j,k\rangle\rangle}\left(\mathcal{J}_{2}\bm{s}_{j}\cdot\bm{s}_{k}+\mathcal{D}_{2}s_{j,z}s_{k,z}\right)$
$+\sum_{\langle\langle\langle j,k\rangle\rangle\rangle}\left(\mathcal{J}_{3}\bm{s}_{j}\cdot\bm{s}_{k}+\mathcal{D}_{3}s_{j,z}s_{k,z}\right)~.$
Here, the unit vector ${\bf\it s}_{j}$ specifies the direction of the spin at lattice site $j$. The first contribution is the Zeeman term, which describes the coupling to the external magnetic field $\bm{h}^{\rm ext}$, while the remaining contributions describe the exchange interaction energy, which we take to be uniform. In particular, $\mathcal{J}_{1}$, $\mathcal{J}_{2}$, and $\mathcal{J}_{3}$ are the exchange coupling constants between nearest (NN), second-nearest (SN), and third-nearest-neighbor (TN) spins, denoted as $\langle\cdot,\cdot\rangle$, $\langle\langle\cdot,\cdot\rangle\rangle\rangle$, $\langle\langle\langle\cdot,\cdot\rangle\rangle\rangle$ respectively. The associated parameters $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$ represent the axial coupling constants, taking into account the fact that the system can have a cylindrical symmetry due to the geometry of the configuration, so they describe the presence of a strong easy-axis or easy-plane anisotropy. This type of uniaxial term explicitly breaks spin-rotational symmetry and can play a crucial role in stabilizing long-range 2D magnetic order [44, 37]. The effective field acting on a given spin at the lattice site $j$
$\bm{h}_{j}=-\frac{\partial\cal H}{\partial\bm{s}_{j}}$ $=\bm{h}^{\rm ext}_{j}-[({\cal J}_{1}\sum\nolimits^{\prime}_{k}+{\cal J}_{2}\sum\nolimits^{\prime\prime}_{k}+{\cal J}_{3}\sum\nolimits^{\prime\prime\prime}_{k})\bm{s}_{k}]$ (2)
$-[({\cal D}_{1}\sum\nolimits^{\prime}_{k}+{\cal D}_{2}\sum\nolimits^{\prime\prime}_{k}+{\cal D}_{3}\sum\nolimits^{\prime\prime\prime}_{k})s_{k,z}]\bm{u}_{z}~,$
where $\bm{u}_{z}$ denotes the unit vector directed along the $z$-direction, and $\sum\nolimits^{\prime}_{k}$, $\sum\nolimits^{\prime\prime}_{k}$, and $\sum\nolimits^{\prime\prime\prime}_{k}$ represent, respectively, the sums over the NN, SN, and TN sites $k$ associated with site $j$. The Hamiltonian given in Eq. (1) has a quite general form, enabling the description of AFM systems with various spin-compensated ordering patterns. Fig. 1 illustrates three possible spin-ordering arrangements for an antiferromagnetic system on a 2D honeycomb lattice, representative of the MPX3 materials [5]. The energetically favored configuration depends on the specific values of the interaction parameters ${\cal J}_{i}$ and ${\cal D}_{i}$ with $i=1,2,3$. In each case, we can decompose the full lattice into two sublattices, denoted $A$ and $B$. Accordingly, we replace the lattice index $j$, which can take $2N$ values, with a pair $(\alpha,p)$, where $\alpha\in\{A,B\}$ and $p$ runs over $N$ values. If the exchange interaction dominates over the anisotropies and external fields, the spatial variations are smooth, and the thermal fluctuations are small, we can apply the macrospin approximation [44]. For each sublattice, we have a collective magnetization, ${\bf\it M}_{\alpha}=N_{\alpha}{\bf\it s}_{\alpha}$, where $N_{\alpha}$ being the spin density on the sublattice $\alpha\in\{A,B\}$, and $N_{A}=N_{B}=N$. Neglecting the external magnetic field and adopting the macrospin approximation, the effective field acting on the spins of sublattice $\alpha$ becomes spatially uniform and can be expressed as
$\bm{h}_{\alpha}=-J\bm{s}_{\bar{\alpha}}-\Delta s_{\bar{\alpha},z}\bm{u}_{z}-J^{\prime}\bm{s}_{\alpha}-\Delta^{\prime}s_{\alpha,z}\bm{u}_{z}~,$ (3)
where $\bar{\alpha}=B$ ($\bar{\alpha}=A$) when $\alpha=A$ ($\alpha=B$). The coefficients $J$ and $J^{\prime}$ ($\Delta$ and $\Delta^{\prime}$) are given as linear combinations of the interaction parameters ${\cal J}_{i}$ (${\cal D}_{i}$), specifically
$J$ $=(Z_{1}-Y_{1}){\cal J}_{1}+(Z_{2}-Y_{2}){\cal J}_{2}+(Z_{3}-Y_{3}){\cal J}_{3}~,$ (4a)
$J^{\prime}$ $=Y_{1}{\cal J}_{1}+Y_{2}{\cal J}_{2}+Y_{3}{\cal J}_{3}~,$ (4b)
$\Delta$ $=(Z_{1}-Y_{1}){\cal D}_{1}+(Z_{2}-Y_{2}){\cal D}_{2}+(Z_{3}-Y_{3}){\cal D}_{3}~,$ (4c)
$\Delta^{\prime}$ $=Y_{1}{\cal D}_{1}+Y_{2}{\cal D}_{2}+Y_{3}{\cal D}_{3}~,$ (4d)
with $Z_{1}$, $Z_{2}$, and $Z_{3}$ denoting, respectively, the number of NN, SN, and TN sites with respect to an arbitrary one (independently of the sublattice), and $Y_{1}$, $Y_{2}$, and $Y_{3}$ representing the corresponding number of NN, SN, and TN sites belonging to the same sublattice. As an illustrative example, we consider the honeycomb lattice shown in Fig 1, where one has $(Z_{1},Z_{2},Z_{3})=(3,6,3)$. Here, the set $(Y_{1},Y_{2},Y_{3})$ depends on the ordering pattern, taking the values $(0,Z_{2},0)$, $(2,2,0)$, and $(1,2,3)$ for the configurations Néel, zig-zag, and stripy, respectively.
Within this classical description, the energetically preferred AFM configuration is obtained by maximizing the magnitude of $\bm{h}_{\alpha}$ under the constraint that it points in the direction opposite to $\bm{s}_{\alpha}$ and parallel to $\bm{s}_{\bar{\alpha}}$. For simplicity, we henceforth restrict our analysis to nearest-neighbor exchange interactions, ${\cal J}_{1}$ and ${\cal D}_{1}$, noting that the formalism developed below can be straightforwardly generalized to include further-neighbor couplings. In particular, we focus on the parameter regime ${\cal J}_{1}>0$ and ${\cal J}_{1}+{\cal D}_{1}>0$, which stabilize an AFM configuration. For a single honeycomb layer, taken as a representative example, this condition clearly stabilizes a Néel configuration. In particular, for $\mathcal{D}_{1}>0$, this model favors alignment along the $z$ direction. Conversely, for $\mathcal{D}_{1}<0$ one obtains a stronger interaction in the $x$-$y$ plane. The spherical symmetry is recovered by taking $\mathcal{D}_{1}=0$. For the sake of clarity, and within the restriction stated above, we explicitly write the Hamiltonian for the two interpenetrating classical spin sublattices (A,BA,B[4, 44],
$\mathcal{H}$ $=-\sum_{\alpha=A,B}\sum_{p}\bm{h}^{\rm ext}_{{\alpha p}}\cdot{\bf\it s}_{\alpha p}+{\mathcal{J}}_{1}\sum_{\expectationvalue{p,q}}{\bf\it s}_{Ap}\cdot{\bf\it s}_{Bq}$ (5)
$+{\mathcal{D}}_{1}\sum_{\expectationvalue{p,q}}s_{Ap,z}s_{Bq,z}~,$
and the effective field acting on a given spin at site $p$ of sublattice $\alpha$ is given by
$\bm{h}_{\alpha p}=-\frac{\partial\cal H}{\partial\bm{s}_{\alpha p}}=\bm{h}^{\rm ext}_{\alpha p}-{\cal J}_{1}\sum\nolimits^{\prime}_{q}\bm{s}_{\bar{\alpha}q}-{\cal D}_{1}\sum\nolimits^{\prime}_{q}s_{\bar{\alpha}q,z}\bm{u}_{z}~.$ (6)
When disregarding the external magnetic field and applying the macrospin approximation, the effective field assumes a homogeneous form and can be rewritten as
@@
where we recall $J=Z_{1}{\cal J}_{1}$ and $\Delta=Z_{1}{\cal D}_{1}$. Before concluding, we briefly outline how corrections due to long-wavelength spatial inhomogeneities can be added to the macroscopic approximation. Starting from Eq. (6), in the absence of the external field, and assuming a generic lattice, we make use of its discrete translational invariance to rewrite the effective field in terms of its Fourier transform as
$\tilde{\bm{h}}_{\alpha}(\bm{P})$ $=\frac{1}{N}\sum_{i}e^{-i\bm{P}\cdot\bm{R}_{p}}\bm{h}_{\alpha p}$ (8)
$=-f(\bm{P})[{\cal J}_{1}\tilde{\bm{s}}_{\alpha}(\bm{P})+{\cal D}_{1}\tilde{s}_{\alpha z}(\bm{P})\bm{u}_{z}]$
where $f(\bm{P})=\sum_{l=1}^{Z_{1}}e^{-i\bm{P}\cdot\bm{d}_{l}}$ denotes the form factor, and $\{\bm{d}_{l}\}_{l=1}^{N}$ represent the NN vectors. As an example, we focus on the honeycomb lattice, for which $\bm{d}_{l}=a_{0}(\sin(l\pi/3),\cos(l\pi/3))^{\rm T}$ [40]. We work in the long-wavelength limit, valid when the characteristic length scale of the inhomogeneities is much larger than the lattice spacing $a_{0}$. In this regime, the form factor can be expanded for small $P$, $P\ll 1/a_{0}$, and approximated as $f(\bm{P})\approx 3(1-a_{0}^{2}P^{2}/4)$.
Under this long-wavelength approximation, we now transform back to real space, so that
$\bm{h}_{\alpha}(\bm{r})=-J\bigg(1+\frac{a_{0}^{2}}{4}\nabla^{2}\bigg)\bm{s}_{\bar{\alpha}}(\bm{r})-\Delta\bigg(1+\frac{a_{0}^{2}}{4}\nabla^{2}\bigg)s_{\bar{\alpha},z}(\bm{r})\bm{u}_{z}~,$ (9)
where we have the replacement $\bm{P}\to i\nabla$, with $\nabla$ representing the nabla operator. It is straightforward to verify that the above expression remains valid for a 2D square lattice, with the form factor $f(\bm{P})=2[\cos(P_{x}a_{0})+\cos(P_{y}a_{0})]$. For a three-dimensional cubic lattice, however, the prefactor is reduced, with $a_{0}^{2}/4$ being replaced by $a_{0}^{2}/6$ [11]. From here on, for the sake of simplicity, we employ the FP formalism within the macrospin approximation. Nevertheless, the discussion presented here could be useful for future extensions that explicitly account for spatial inhomogeneities.
@@ II.2 Fokker-Planck formalism
Here, we derive the dynamics equation for the probability distribution of the magnetization configuration of an AFM, starting from a stochastic version of the LLG equation that includes Langevin fields to incorporate the effects of thermal fluctuations  [11]. For each of the two sublattices constituting the AFM, the stochastic LLG equations are expressed as
@@
where $\dot{{\bf\it s}}_{\alpha p}=\partialderivative{{\bf\it s}_{\alpha p}}{t}$, $\alpha\in\{A,B\}$, $\bm{h}_{\alpha p}$ is the effective magnetic field associated with the Hamiltonian of the spin defined in Eq. (6), $\bm{\zeta}_{\alpha p}$ is a Langevin field that is postulated to be uncorrelated at different sites. Note that we are working in units such that the lattice-independent [4], gyromagnetic ratio $\mu=1$, and $\hbar=1$. Both $\Lambda$ and $\Lambda^{\prime}$ are two diagonal matrices
@@
where $\lambda,\lambda+\delta\lambda\ll 1$ are the intra-sublattice Gilbert damping factors [16, 17], and $\lambda^{\prime},\lambda^{\prime}+\delta\lambda^{\prime}\ll 1$ are the inter-sublattice Gilbert damping factors [24, 23]. Inclusion of parameters $\delta\lambda$ and $\delta\lambda^{\prime}$ allows uniaxial symmetry in damping mechanisms, following the cylindrical symmetry of the spin-spin interactions.
@@
Following Ref. [12], Eq. (12) can be rewritten in the form of the (componentwise) Langevin equation
@@
Langevin fields $\bm{\zeta}_{\alpha}$ are considered stochastic processes [46, 12, 11]. Assuming that they are Gaussian and that they satisfy the following statistical properties
@@
Moreover, we underline that the condition $D_{\mathrm{T},j}\geq 0$ leads to a constraint on the anisotropy $\delta D_{\mathrm{T}}\geq-D_{\mathrm{T}}$. Using the quantities in Eqs. (14) and (15), it is possible to obtain the coefficients of the Kramers-Moyal expansion, which describes the evolution of the probability distribution function (PDF) for a generic Markov process [27, 35, 12] (see Appendix A for further details). One obtains the FP equation [12], namely the equation governing the dynamics for the PDF, $\mathcal{P}({\bf\it\mathcal{S}},t)$,